OpenStudy (anonymous):

WILL AWARD MEDAL! George has found a relationship between the number of trees in an orchard and the number of species of birds found in the orchard. The table below shows the data collected by George: Number of trees (x) 100 200 300 400 500 600 700 800 900 Number of birds (y) 51 101 151 201 251 301 350 400 450 Part A: What would most likely be the number of birds in the orchard if there are 284 trees? Part B: Predict a possible correlation coefficient for the data in the table and explain why you think your prediction is a good value for the data.

OpenStudy (anonymous):

Part C: George says that if water sources are provided in the orchard, the bird population would increase. Is this an example of correlation or causation? Justify your answer.

OpenStudy (anonymous):

@marissalovescats

OpenStudy (anonymous):

@austinisawesome12

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

let em see

OpenStudy (anonymous):

@Clalgee

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@imer

OpenStudy (kirbykirby):

DO you know the regression line equation?

OpenStudy (imer):

Answer for part A should be "143" 300 Trees = 151 Birds 200 Trees = 101 Birds 300-200=151-101 Difference: 100 Trees = 50 Birds Twice as many trees as birds. Therefore if we follow the pattern; 300-284=x Difference: 16 Trees = 8 Birds

OpenStudy (kirbykirby):

The regression line is \[y=\hat{\beta_0}+\hat{\beta_1}x \] where \[\hat{\beta_0}=\bar{y}-\hat{\beta_1}\bar{x} \] and \[\hat{\beta_1}=\large \frac{\sum_{i=1}^n(x_i-\bar{x})(y_i-\bar{y})}{\sum_{i=1}^n(x_i-\bar{x})^2} \]

OpenStudy (kirbykirby):

Once you find \(\hat{\beta_0}\) and \(\hat{\beta_1}\), go back to the line equation \(y=\hat{\beta_0}+\hat{\beta_1}x\) and plug in \(x=284\)

OpenStudy (kirbykirby):

It should be close though to 143 @imer said because the data is almost perfectly linear (as in question b), the correlation coefficient is veryyyy close to 1, and is probably around 0.999 or smtg like that)

OpenStudy (kirbykirby):

You can probably just use software (R, SPSS, even Wolfram I think can do it) though to fit the line using your data, It's kinda long to do that by hand

OpenStudy (imer):

@kirbykirby The question is not sophisticated as you can see the word "most likely" in part "A" and since the values are discrete, I am pretty sure the question does not demand knowledge about regression line.

OpenStudy (kirbykirby):

Hm ya as I was answering myself, I think I realized it's not such a sophisticated question as you mentioned, although that realization came from question b) to me as it said to predict the value.. and certainly any program would just spit it out haha. Sorry for the confusion on my part :(

OpenStudy (anonymous):

so for part b our predicted correlation coefficient value is 1? and what about part C?

OpenStudy (anonymous):

you guys have been an incredible help by the way

OpenStudy (kirbykirby):

Probably correlation. Proving causation is actually quite difficult. Also, there is very little information in the question that says what George has done in the study to show this. But most likely, he did an observational study (you don't control the experiment yourself, meaning he can't control the number of birds visiting the orchard). Typically you can try to reach causation by doing an randomized experimental study. You can assign things to certain groups (treatment and control) by randomizing them in order to "cancel" out any possible uncontrollable outside effects ("lurking variables"). I'm not sure how much of this you have study, but the general saying is "correlation does not imply causation". So while George may notice an increase in the number of birds with the addition of a water source (a positive correlation), it does not imply causation.

OpenStudy (kirbykirby):

You can do some kind of proving causation in an observational study, but the criteria to do so are very rigorous.